1 December 2009 Lab Meeting
- Oakridge deadline is 1 Feb 2010.
- Need new lab meeting time for Spring.
- Shichu Huang will be joining us in Feb.
‡ Flu R01:Integration
† HDA (Lead: Jaephil, Team:Sonali)
- Start planning R01 for Submission on 2/5/10. MM, JD, CMK.
- Dry reagent storage for RT and HDA - COP test well array provided to Sonali
- 2nd HDA design - include RT, dry reagent, reduced evap loss, and parallel filling, etc
- 2nd HDA meeting - Dec 10, 12-2pm
- Paper resubmitted to BMMD 11/30/09.
- "prototyping - scratch" goes to "Special Issue of the Robotics and Computer-Integrated Manufacturing Journal" by DEC 31.
† Sample Concentration (Lead: Jane, Team: Jaephil)
- New design sent to make Rubber mold: outline octagon problem. Revised design sent yesterday.
- Jane will train Cassidy once more on Wednesday and Thursday this week.
- MM 040808 A/PR/8/34 flu stock with PFU = 2X10^4 available to Jane to train Cassidy in ultracentrifuging all tubes
- of virus, pour out sucrose,pipette remaining with 20uL pipette tip and pool for plaque assay for supplying virus
- Jane working on the cell lysate control. - have many tubes now, done
- Set up COMSOL and StarCD, look for governing equations for simulation of evaporation. - ongoing
† SPE Column Optimization for DNA/RNA. (Lead: Sonali, Team: Hussam, Jessie, Brendan)
- pH the sample. Optimum 6.6. Make sure that we are there. 3M GUSCN stands at pH 4.2 very acidic (some NaOH)(11/24)
- "Measure again 1:1 with DEPC water, 1:1 with PBS" (11/24)
- 3M GuSCN in water (suspended from 6M to 3M in DEPC) + sample(DEPC) [1:1] pH~4.08
- 3M GuSCN in PBS (suspended from 6M to 3M in 2X PBS) + sample(DEPC) (1:1) pH~6.43
- DEPC water has pH=4.19
- Silica nanoparticles from Fluka. 15nm "order"
- Still working on the 150nm particles 1X experiments with lambda. Problem with bonding, had to optimize Jaephil's technique.
for our chip, now it's better.
- SEM pictures of the 150nm Silica nanoparticles.
- Dispersion of particles? Tween or another dispersant?
- PLanning net silica per channel experiments. with 700nm, 3x possible. Deliver around 11/5/09.
- 150nm and 700nm silica with 0.5mm (3.6uL SPE) channels.
- 150nm Silica did not improve Ct on average, but one channel with 700nm silica performed exceptionally well, which brought the average up.
- "relationship between PFU/ml and copy # of RNA. Do this via a serial dilution of Virus, put through Qiagen kit. Assuming that Qiagen is 100% efficient."
- "Jessie get trained for real on SEM.*"
† PCR - CMI (Lead: Qingqing)
- QQ will work on the initial integration steps of SPE + RT (reservoir)+ PCR.
- SPE+RT+PCR chip is designed.
- mold has been made.
- toxin A.
HDA dilution experiments for toxin A has been tested in tube.
the products has been tested in Bio-analyzer. 1pg is the limitation for the template, in order to get a detectable result.
- PCR2 Paper formatted for LOAC this week.
need to get back from Cathie
- CMK: PCR 1 draft. Analytical Chem. MCK is running more simulations.
‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )
- Sonali to train Lisa on SPE.
- They don't have syringe pumps and have not asked for training yet. I offered training here which was declined.
- QQ to run test PCR on chip with genomic DNA and Toxin B primers.
‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)
- Meeting 9am Monday (every other week)
- "ASB will give action items."
‡ Agilent Automated Sample Preparation (Lead: Alex)
- So far tested on SNAP2:
- B. subtilis gDNA
- B. subtilis cells, high Cts for the range tested.
- Planned tests SNAP2:
- B. subtilis cells (higher concentrations)
- MDCK gDNA
- HotDog Sample
- In Progress: Machining of pipette-tip-holder to run Nano-C tests
- First lysis test with B.Subtillis cells planned for end of this week.
- Paper on Yeast: Extended Literature Review (12/11)
‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)
- New metal piece - First week of Dec.
- Evap paper - short talk with Larry before COBRA meeting
‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)
- Cathie will submit paper inquiry.
‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)
- IRB approved.
- BU IRB was approved.
‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME))
- Submitted Gantt Document to PATH 11/25.
- Senior Project proposal presentations this Friday, Dec. 4.
- Frank: PVA experiments to begin shortly.
- Sean: Design has been submitted to Alex and awaiting the product. I can then get back to working on the optimization of the straws, ie the best monolith composition for extracting DNA/RNA onto the straws
- Mark to begin initial DNA testing with shorter time periods (1,3 days, etc.) to be completed this(?) week. New straw fixture to be set up in old fume hood. Fume hood to be sterilized for use only with straw machine.
‡IIH Senior Project.
- Try Dow Chemical Website. FOOD safe, FDA grade.
- Issue finding double-stick material
- currently inquiring about a PEEK tape and a PC film with adhesive
- Non-double stick material
- fluorinated ethylene propylene, PEEK and PC are available from McMaster Carr (no acrylic)
- Appropriate chip design software?